Someoneee:

15x^3 - 12x^2 + 20x - 16

Someoneee:

snowflake0531:

Let's do it by group factoring 15x^3 - 12x^2 + 20x - 16 Let's first factor the first part, 15x^3 -12x^2 What is the GCM?

Someoneee:

1?

Someoneee:

wait nvm, 3

snowflake0531:

there's also the x^2

Someoneee:

x^2

Someoneee:

yes

snowflake0531:

lol

snowflake0531:

15x^3 -12x^2 3x^2(5x-4) yes?

Someoneee:

yeah

snowflake0531:

and then the second half 20x-16 factor that

Someoneee:

4

snowflake0531:

So 4(4x-4) yea?

Someoneee:

yep

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @snowflake0531 So 4(4x-4) yea? \(\color{#0cbb34}{\text{End of Quote}}\) 5x not 4x

snowflake0531:

ye, 5x, oops So, we have 3x^2(5x-4) + 4(5x-4) Do the opposite of how you would usually distribute things Remember ab+ac = a(b+C)

snowflake0531:

Imagine the 5x-4 as just one letter, it'll make it easier to think about

Someoneee:

So, then the answer would be (3x^2 - 4)(5x-4)? right?

snowflake0531:

Huh? Let's first imagine the 5x-4 as y So, it would be 3x^2 y + 4y Do the opposite of distributive property for ^^^ that

snowflake0531:

Like, a(b+c)=ab+ac ab+ac=a(b+c)

Someoneee:

The answer would be (3x^2 + 4)(5x - 4). My mistake.

snowflake0531:

Yepppp